If is normally distributed with mean 4 and variance 4, find the probability of 1.202<X<83180000. (Given that log10 1202 =3.08,log10 8318=3.92)
1
Expert's answer
2020-04-19T15:08:47-0400

"P=P(1.202<X<83180000)" is similar in log normal as,


"P=P(\log_{10}1.202<\log_{10}X<\log_{10}83180000)\\
 P=P(\log_{10}\frac{1202}{1000}<\log_{10}X<\log_{10}(8318*10^4))\\
 P=P(3.08-3<\log_{10}X<3.92+4)\\
 P=P(0.08<\log_{10}X<7.92)\\"


converting to normal form using,

"z=\frac{x-mean}{standard\ deviation}=\frac{x-4}{2}"


"P=P(-1.96<z<1.96)\\
P=P(-1.96<z<0)+ P(0<z<1.96)\\"

using calculator or table,

"P=0.475+0.475\\
\bold{P=0.95}"